OpenStudy (anonymous):

Simplify:10 a squared over 7 b squared divided by 25 a cubed over 14 b cubed

OpenStudy (amistre64):

simplify how? by writing it in mathematical notation?

OpenStudy (anonymous):

by distributive property

OpenStudy (amistre64):

i dont think i can parse a distribution from it

OpenStudy (anonymous):

thats what i dont get because i dont know where to start with the distribution part of the equation

OpenStudy (amistre64):

10a^2/7b^2 ----------- 25a^3/14b^3 2 4b ----- = --- 5a/2b 5a

OpenStudy (anonymous):

can you explain to me where to began please

OpenStudy (amistre64):

well, i didnt distribute ... so i have no idea where to begin :)

OpenStudy (anonymous):

what did you use to solve it is probably way easier

OpenStudy (amistre64):

i multiplied by 1 \[\large \frac{\frac{10a^2}{7b^2}}{\frac{25a^3}{14b^3}}*\frac{\frac{14b^3}{25a^3}}{\frac{14b^3}{25a^3}}\] \[\large \frac{\frac{10a^2}{7b^2}*\frac{14b^3}{25a^3}}{1}\] \[\frac{10a^2}{7b^2}*\frac{14b^3}{25a^3}\] \[\frac{10a^214b^3}{25a^37b^2}\] \[\frac{10\cancel{a^2}14b}{25a7\cancel{b^2}}\] \[\frac{\cancel{10}2\ \cancel{14}2b}{\cancel{25}5a\cancel{7}}\] \[\frac{4b}{5a}\]

OpenStudy (anonymous):

oh okay thank you greatly appreciate it :)